A man climbs a ladder carrying a hammer of mass 0.5kg.If the ladder makes an angle of 30 with the vertical and is 3m long, How much work does the man do on the hammer?
1
Expert's answer
2020-09-14T10:20:18-0400


The work done against the gravitational force is:


"W = mgh"

where "m = 0.5 kg" is the mass of the hummer, "g = 9.81 m\/s^2" is gravitational acceleration and "h" is the difference between the final and initial levels. From the figure above:


"h = 3\cos{30\degree} = \dfrac{3\sqrt3}{2} m"

Thus, the work will be:

"W = 0.5\cdot 9.81\cdot \dfrac{3\sqrt3}{2}\approx12.74 J"

Answer. 12.74 J.